# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance

### v1

@@ IX-E Proof of Lemma 4
For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}),\beta\in[0,\gamma_{a}]$, according to claim 1 in Lemma 4, one has $\mathbf{u}(\xi)\in(\mathbf{T}_{(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})}(\mathbf{x}))^{\circ}$. This implies that when the control input corresponds to the obstacle-avoidance mode, it will steer the state $\mathbf{x}$ towards the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
@@
$\mathbf{u}(\xi)=-\kappa_{r}{\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}},\kappa_{r}>0.$ (49)
For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}),\beta\in[\gamma_{s},\gamma]$, according to claim 1 in Lemma 4, one has $\mathbf{u}(\xi)\in(\mathbf{T}_{(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})}(\mathbf{x}))^{\circ}$. This implies that when the control input corresponds to the obstacle-avoidance mode, it will steer the state $\mathbf{x}$ towards the interior of the set $(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
Finally, we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi})$. When the control law operates in the obstacle-avoidance mode and the state $\mathbf{x}\in\mathcal{S}_{2}$, according to (IV-A), one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a}(t_{1},j_{1}+1)),\kappa_{r}>0.$ According to (12), it can be confirmed that $\mathbf{v}(\mathbf{x},\mathbf{a}(t_{1},j_{1}+1))^{\intercal}\mathbf{R}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{P}(\mathbf{a}(t_{1},j_{1}+1))\mathbf{x}_{\pi}>0$, when the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2},$ and the proof is complete.

### v2

@@ IX-E Proof of Lemma 4
If we show that for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$, then it implies that the control input vector (49) steers $\mathbf{x}$ to the interior of the set $\mathcal{N}_{\gamma-\beta}(\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i}))$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 4, ensures that the control input vector (49) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$. To proceed with the proof we require the following fact:
Fact 1: Let us consider a hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$, where $\mathbf{p}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for some $i\in\mathbb{I}$ and $\mathbf{q}=\mathbf{A}(\mathbf{p})$, where the mapping $\mathbf{A}$ is defined in (29). Then, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}>0$.
Proof.
This proof is by contradiction. First, note that for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{q})$ such that $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that that plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [19, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.
However, since $\mathbf{q}=\mathbf{A}(\mathbf{p})$, according to (29), one has $\mathbf{p}_{\pi}\in\mathcal{P}(\mathbf{p},\mathbf{q})$. Therefore, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\subset\mathcal{P}(\mathbf{p},\mathbf{q})$. As a result, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$. This implies that $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})\neq\emptyset$, which is a contradiction. ∎
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, where $\beta\in[0,\gamma_{a}]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. Therefore, as discussed earlier, the control input vector (49) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
@@
$\mathbf{u}(\xi)=-\kappa_{r}{\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}},\kappa_{r}>0.$ (50)
Let $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$ for some $\beta\in[\gamma_{s},\gamma]$. We know that for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, the tangent cone to the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$ is given by
$\mathbf{T}_{\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})}(\mathbf{x})=\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi}).$
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, where $\beta\in[\gamma_{s},\gamma]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. This implies that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 4, ensures that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
Finally, we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2}$, the control vector $\mathbf{u}(\xi)$ belongs to the open positive half-space $\mathcal{P}_{>}(\mathbf{0},\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi})$. When the control law operates in the obstacle-avoidance mode and the state $\mathbf{x}\in\mathcal{S}_{2}$, according to (12), one has $\mathbf{u}(\xi)=\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{a}),\kappa_{r}>0.$ Note that for all $\mathbf{x}\in\mathcal{S}_{2}$, one has $\eta(\mathbf{x})\in(-1,1)$. Therefore, for every $\mathbf{x}\in\mathcal{S}_{2}$, the vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ can be expressed as a linear combination of the vectors $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$ and $\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$ given by
$\mathbf{v}(\mathbf{x},\mathbf{a})=k_{1}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}+k_{2}\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi},$ (51)
where $k_{1}\in\mathbb{R}$ and $k_{2}>0$. Additionally, according to Fact 1, for all $\mathbf{x}\in\mathcal{S}_{2}$, one has $\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\neq\mathbf{0}.$ As a result, it can be confirmed that $\mathbf{v}(\mathbf{x},\mathbf{a})^{\intercal}\mathbf{R}(\mathbf{a})\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$, when the state $\mathbf{x}$ belongs to the set $\mathcal{S}_{2},$ and the proof is complete.